Hypothesis Testing

According to the Journal of Counseling Psychology, forbearance is a coping strategy of collectivist society with 150 Chinese respondents scoring an average of 35 with a standard deviation of 7 in a ten item forbearance scale. You want to test if Europeans use the same coping strategy andyou found that they scored an average of 25 in the same forbearance scale in a thesis. Discuss your finding findings at α 0.05 in traditional and p-value method.
1
Expert's answer
2020-05-18T18:42:41-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=35"

"H_1:\mu\not=35"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.05," and the critical value for a two-tailed test is "z_c=1.96."

The rejection region for this two-tailed test is "R=\{z:|z|>1.96\}."

The z-statistic is computed as follows:


"z={\bar{X}-\mu\over \sigma/\sqrt{n}}={25-35\over 7/\sqrt{150}}\approx-17.496"

Since it is observed that "|z|=17.496>z_c=1.96," it is then concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu"  is different than 35, at the 0.05 significance level.

Using the P-value approach: The p-value is "p=0," and since "p=0<0.5," it is concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is different than 35, at the 0.05 significance level.